OpenStudy (gojanij):

A golden rectangle has side lengths in the ratio of about 1 : 1.618. If the long side of a golden rectangle is 35 cm, what is its area? Round your answer to the nearest tenth.

OpenStudy (whpalmer4):

In a golden rectangle, the side lengths are in the ratio \[\text{short:long} \approx 1:1.618\]or written as fractions \[\frac{\text{short}}{\text{long}} = \frac{1}{1.618}\]You know what the long side is, so plug it into the equation, cross-multiply and solve for the short side.

OpenStudy (gojanij):

21.632= shortside?

OpenStudy (whpalmer4):

The exact value is \[1:\frac{1+\sqrt{5}}{2}\]which has an unending decimal representation. One of the many nifty things about it is that if you take the reciprocal of \(\phi = 1.618...\) you get the same string of digits after the decimal point! \[1.6180339887498948482045868343656381177203091798057628621354486227\]\[0.6180339887498948482045868343656381177203091798057628621354486227\] You are close: correct manipulation, but haven't followed the directions on rounding your answer...

OpenStudy (whpalmer4):

There's a fun book on the Golden Ratio: http://www.amazon.com/The-Golden-Ratio-Worlds-Astonishing/dp/1469286092

OpenStudy (gojanij):

A. 2450 cmlook at the picture B. 757.1 cmlook at the picture C. 612.5 cmlook at the picture D. 1982.1 cmlook at the picture

OpenStudy (gojanij):

ignore the look at the picture part but the answers are to the second power

OpenStudy (whpalmer4):

okay, so you have the long side = 35 and the short side = 21.6 and you need the area. area of a rectangle is given by \[A = l*w\]

OpenStudy (gojanij):

i got 756

OpenStudy (whpalmer4):

35*21.632 = 757.12 Don't round until you reach the final answer to prevent loss of accuracy such as you experienced.

OpenStudy (gojanij):

oh ok thank you very much!

OpenStudy (whpalmer4):

You're welcome!